OpenStudy (anonymous):

Please Help! I always give medals. 5. McSaw’s Ice Cream makes muffins that cost $.60 each. Past experience shows that 10% of the muffins will become stale and have to be discarded. Assume McSaw wants a 55% markup based on cost and produces 100 muffins. What should each muffin sell for? Round to nearest cent.

OpenStudy (anonymous):

\[100 - 10 = 90 \times .6 = $54 \times 1.55 = 83.7 \div 90 = .93 \]

OpenStudy (anonymous):

So your answer is 0.93 cents a muffin

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

No problem.